Calculate molar concentration of a sodium hydroxide solution,if 25cm³ of it requires 20.5cm³ of 0.103 mol/dm³ solution of tetraoxosulphate(VI) acid for neutralization.what is the concentration of the alkaline solution in g/dm³


1
Expert's answer
2021-03-23T07:25:37-0400

Molar Mass of sodium hydroxide= 39.997 g/mol

25cm3 = 25 g

= 25/39.996 = 0.625048 mol

Molar Mass of tetraoxosulphate (VI) H2SO4 = (2+32+64) = 98g/mol

0.103 × 98 = 10.094g

20.5g of 10.0094 dm3= 20.5/10.904 = 1.880044

= 0.625048/ (0.025 + 1.880044)

= 0.328101mol/L